OpenStudy (anonymous):

19. Tell whether the lines for each pair of equations are parallel, perpendicular, or neither. y=-7/8x-1 32x - 28y=-36 (1 point) A.parallel B.perpendicular C.neither

OpenStudy (mertsj):

Are the slopes the same? If so they are parallel Are the slopes negative reciprocals of one another? If so they are perpendicular No to both of those questions? Then they are neither

OpenStudy (anonymous):

@alquana, get them both in the form y =mx +b. See for yourself!

OpenStudy (anonymous):

ok i did but i'm just checking to see if i be having the correct answer!

OpenStudy (mertsj):

What is your answer?

OpenStudy (anonymous):

neither

OpenStudy (anonymous):

Did you solve for y with this equation? 32x - 28y=-36

OpenStudy (mertsj):

The slope of the first one is 7/8 The slope of the second one is -8/7 -8/7 is the negative reciprocal of 7/8 so therefore they are??????

OpenStudy (mertsj):

Whoops. I got the signs backwards. The slope of the first one is -7/8 The slope of the second one is 8/7 The conclusion is the same.

OpenStudy (anonymous):

It's okay. So the first slope is -7/8 and 8/7 Hey Alquana, go onto google and graph: y= -7x/8 and: y=8x/7 See what they both look like.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Does that look par, perp, or neither?

OpenStudy (anonymous):

to me neither

OpenStudy (anonymous):

http://www.youtube.com/watch?v=0671cRNjeKI